OpenStudy (anonymous):

x^2/x+5 = 25/x+5 solve the equation

OpenStudy (lgbasallote):

\[\huge \frac{x^2}{x+5} = \frac{25}{x+5}\] multiply both sides by x+ 5\[\huge \implies x^2 = 25\] take the square root of both sides \[\huge \implies x = \pm \sqrt{25}\] \[\huge \implies x = \pm 5\] does that help?

OpenStudy (anonymous):

yeap

OpenStudy (lgbasallote):

wonderful